# BSTcontrol

### v2

@@ 3.3 WLAM for Semantic Communications
SemCom is a key technology for future 6G systems, promising superior communication efficiency by directly optimizing information transfer at the semantic level [133, 134, 135, 136, 137]. Large AI models are pivotal for realizing SemCom, providing the essential capability to extract and process semantic information from diverse data modalities like images, text, audio, video, and knowledge graphs, tailored to specific tasks such as classification and reconstruction. Key challenges for large AI models in SemCom include accurately identifying and precisely transmitting the desired meaning within communication systems, and ensuring robustness against channel variations through adaptive designs and re-transmission mechanisms. Leveraging the power of large AI models to address these challenges is anticipated to unlock substantial performance gains in semantic communication efficiency and reliability. The communication flow chart is demonstrated in Fig. 9.
@@ 3.3.1 Semantic Communications Preliminaries
Understanding the theoretical underpinnings of SemCom is crucial for analyzing its performance limits and guiding system design [135, 134, 138, 139, 140]. Early work reintroduced the concept of semantic entropy, positing that semantic units exhibit logical connections rather than being random [133]. Building on this, the idea of utilizing shared knowledge bases between communicating parties to exchange semantic information highlighted its potential [141]. Subsequent research provided more formal definitions, such as considering multiple source units as equivalent if they share the same meaning (synonymous mapping), leading to generalized definitions of semantic entropy and insights into the fundamental limits of SemCom, particularly for many-to-one source scenarios [142]. It is recognized that this synonymous mapping, where different expressions convey identical meaning (e.g., She appeared happy vs. She appeared joyful), is a key factor enabling SemCom to potentially outperform traditional bit-level communication. These developments have contributed to establishing a systematic framework for semantic information theory [138].
@@ 3.3.2 Modality/Task Tailored SemCom
In text domain, the authors in [143] first proposed deep learning based joint source and channel coding (DeepJSCC) for text transmission. They developed an recurrent neural network (RNN)-based network to directly map text to channel symbols, bypassing the separate steps of data compression and interference resistance by source coding and channel coding. This method demonstrated a lower word error rate (WER) compared to conventional methods by effectively leveraging the structural feature of text. Building on the initial idea of using AI for feature extraction, a comprehensive transmission framework for text called DeepSC was proposed in [144]. DeepSC employs the advanced transformer architecture for text feature extraction and reconstruction. Training the transformer model in an end-to-end manner significantly improved performance. Additionally, they introduced a new performance metric called bilingual evaluation understudy (BLEU), which measures the similarity between two sentences in the feature space. Unlike WER, BLEU aligns more closely with human-like semantic definitions. To address the complexity concerns, a lite distributed semantic communication system called L-DeepSC was proposed in [145], where network pruning and quantization are adopted to reduce the computation complexity of network part. Recognizing the importance of semantic-aware metric, the authors in [146] further proposed a metric of semantic similarity (MSS) that compares two sentence through their graph similarity. Based on which, they optimizes the system performance by reinforcement learning with MSS as reward.
@@
In image domain, substantial research efforts have been devoted to improving SemCom. Initially, the authors in [147] proposed the CNN models for image transmission, employing convolutional layers for encoding and transpose convolutional layers for decoding. Their experiments demonstrated the potential of DeepJSCC in image transmission by significantly outperforming the conventional schemes (i.e., JPEG+LDPC) in the low SNR regime. This approach was further refined in [148, 149]. With the advent of powerful vision transformers, researchers began exploring transformer-based DeepJSCC networks. A vanilla transformer model for DeepJSCC was proposed in [150]. This model was subsequently improved in [151, 32] by replacing the original transformer block with swin transformer block. Recently, the authors in [152] introduced the mamba model for DeepJSCC, achieving better performance than transformer-based JSCC.
@@
In the video domain, the authors in [153] were the first to consider end-to-end DeepJSCC for video transmission. They classified video frames into key frames and non-key frames. Key frames are transmitted using the image-style JSCC model proposed in [154], while non-key frames are transmitted by encoding the residual information relative to the key frames. Additionally, they used reinforcement learning to achieve resource allocation at the frame level based on motion strength, demonstrating numerical results that overcome the cliff-effect present in conventional transmission schemes. This approach was further improved in [155]. Subsequently, the authors in [156] proposed using the last frame as a condition for the encoding and decoding process. This paradigm allows for the full exploitation of temporal dependence and enables resource allocation for each feature embedding. Their experiments showed better performance than the conventional separate source and channel coding schemes. More recently, the authors in [157] considered video transmission in static scenarios, representing the movement of objects as a graph. This approach achieved better performance than conventional schemes.
@@
In audio domain, similarly, the structural feature of speech signal can be leveraged to reduce the transmission overhead or combat the channel noise. The authors in [158] first proposed a deep learning based SemCom system for speech signals, called DeepSC-S. They utilised the attention mechanism to identify the important information and transmit them with larger power. This approach demonstrated their effectiveness by outperforming the conventional audio transmission scheme under the common speech signals metrics. Realizing the correlation between speech and text, DeepSC-S was further extended to speech synthesis [159]. In the speech to speech task, the authors in [160] considered two scenarios, that is, speech recognition and reconstruction. For the recognition task, the semantic-relevant information was extracted and transmitted to the receiver. For the reconstruction task, some additional information that is irrelevant to the semantic information is also transmitted for consistent speech synthesis.
@@
In addition to the common modalities in media form, SemCom can also be applied to knowledge graphs, which are regarded as a standard and universal modality. Knowledge graphs can effectively express the attributes and their connections, thereby denoting semantic information [161, 162]. The authors in [163] first considered adopting knowledge graphs for wireless image transmission. They extracted the knowledge graph from the image to be transmitted selectively, and transmitted the semantic information based on its importance and the available resource blocks. Leveraging on the feature of knowledge graph, they proposed a multimodal performance metric called image-to-graph semantic similarity (ISS) to measure the transmission accuracy of semantic information. Additionally, a multi-agent RL algorithm was proposed to minimize the sum of the average transmission accuracy while satisfying ISS requirement. Then, the transmission of knowledge graph in multi-user scenario was further considered in [164, 165, 166].
@@ 3.3.3 Generative AI Models for SemCom
DeepJSCC has marked significant progress in SemCom, primarily by minimizing the distortion between the reconstructed data and the original source. However, just focusing on low distortion does not always guarantee optimal semantic performance or perceptual quality, reflecting the inherent rate-distortion-perception tradeoff [101, 167]. To bridge this gap, researchers are increasingly integrating generative AI models into SemCom systems. These models, capable of learning complex data distributions and generating high-fidelity samples, offer powerful tools to enhance semantic fidelity and perceptual quality beyond traditional distortion metrics. Two main categories of generative models are being explored below.
@@
This line of research focuses on using established generative techniques like GANs and diffusion models to improve the quality of data reconstructed by DeepJSCC, particularly for perceptual tasks. One strategy involves employing generative models as post-processing modules. After DeepJSCC reconstructs the data, a generative model refines the output to look more realistic, such as using StyleGAN for face images [101] or diffusion models for general image restoration [167]. The primary challenge in wireless communications is accurately modeling the complex distortion introduced by both the DeepJSCC model and the noisy wireless channel [168]. An alternative approach leverages the inherent denoising capabilities of diffusion models more directly, using them either for preprocessing or as integrated denoisers at the receiver to combat channel impairments [102]. Recognizing the computational and bandwidth demands of these models, techniques like integrating compression networks [169] and knowledge distillation [170] have been explored to improve efficiency.
@@
Leveraging the unique capabilities of LLMs opens fundamentally new avenues for SemCom, shifting focus towards deep semantic understanding and knowledge utilization. LLMs can serve as expansive knowledge bases, enabling highly efficient communication where the transmitter sends minimal, semantically crucial information, and the receiver-side LLM reconstructs the full meaning using its stored knowledge and contextual reasoning [103, 171]. Furthermore, the training objective of LLMs, related to entropy minimization [172], makes them inherently effective source coders capable of semantic compression. The extension of LLMs to multimodal data allows for their integration into multimodal SemCom systems, enhancing information exchange across diverse formats. Finally, there is a synergistic interplay where communication can update the knowledge of LLM, and the refined knowledge, in turn, improves communication efficiency and context-awareness.
@@ 3.4 Wireless Agents
Large AI models like LLMs are transitioning from tools for specific tasks like text generation or analysis into more autonomous entities, known as agents capable of planning reasoning and interacting with environments to achieve complex goals [106, 173, 174]. This agentic paradigm leverages the advanced understanding reasoning and generation capabilities inherent in LLMs applying them to orchestrate complex systems like future wireless networks. In 6G communications LLMs as agents promise to manage network complexity automate operations and enable novel intelligent services by perceiving the network state making informed decisions and interacting with network functions through defined interfaces or APIs [175, 176], as illustrated in Fig. 10. This shift requires specific system designs model training methodologies and interaction frameworks to harness their potential effectively and reliably within the telecommunications domain. Implementing these complex agentic systems is further facilitated by the development of dedicated platforms, such as the open-source OpenManus [177] framework designed for building general AI agents.
@@ 3.4.1 Telecom Knowledge Agent
LLMs are increasingly being adapted to function as specialized telecom knowledge agents, capable of navigating the complexities of the telecommunications domain. TelecomGPT offers a representative framework for this adaptation, proposing a comprehensive pipeline to transform general purpose LLMs into telecom specific variants [104]. This process utilizes collected telecom data for continual pretraining, instruction dataset construction for fine tuning, and preference data for alignment tuning. The TelecomGPT framework also introduces novel benchmarks designed to evaluate critical model capabilities, including mathematical modeling, question answering, and code generation specific to the telecom context.
Specialized models enhance information retrieval from technical documents, a crucial task for telecom professionals. TeleRoBERTa, for example, provides an extractive question answering model highly effective at referencing 3GPP documents [178]. By adapting the RoBERTa base model and fine tuning it on telecom data, TeleRoBERTa delivers performance comparable to much larger foundation models on benchmarks like TeleQuAD, showcasing the efficiency gains possible through domain specific training even with fewer parameters. Complementing this, TelecomRAG leverages retrieval augmented generation to create a telecommunication standards assistant [77]. This system focuses on providing accurate, technically detailed, and verifiable responses, directly addressing the limitations of generic models regarding precision and source traceability when dealing with standard specifications.
The development of domain specific models is further supported by initiatives like the Tele-LLMs series, which provides open source LLMs adapted for telecommunications [179]. These models, enhanced through continual pretraining on telecom related data, demonstrate improved performance on domain relevant tasks compared to their general purpose predecessors. Together, these frameworks and specialized models highlight significant advancements in tailoring LLMs to serve as effective and reliable knowledge agents within the telecommunications field, promising greater efficiency and accessibility for professionals navigating complex technical information.
@@ 3.4.2 Code Generation
General applications include translating natural language descriptions into executable code, providing context aware code completion suggestions during development, and performing automatic program repair to correct bugs in existing software [180]. Beyond these general uses, studies have shown the applicability of these models in more specialized areas relevant to communications engineering. For instance, LLMs have been successfully employed to generate complex hardware description language code, such as Verilog, for implementing advanced wireless communication algorithms like the fast Fourier transform on field programmable gate arrays. Techniques including in context learning and chain of thought prompting were utilized to navigate the intricacies of hardware specific requirements like subtask scheduling and multi step reasoning, which are less common challenges in standard software code generation [105]. Another targeted application involves the automated generation of test scripts essential for validating telecom software systems. A proposed two stage framework first uses generative models to create synthetic yet realistic test input data based on historical network performance, and then employs a LLM to generate the actual test script code using this synthetic data combined with natural language test descriptions [132]. This addresses the tedious nature of manual test creation and helps cover diverse scenarios, crucial for complex telecom environments like O-RAN.
@@
While the application potential is vast, the evaluation of code generated by LLMs presents ongoing challenges. Much current evaluation work focuses primarily on functional correctness, often assessed through pass rates on standardized programming benchmarks, and on identifying security vulnerabilities [180, 181]. However, research indicates a gap between the advancement of generation capabilities and the comprehensiveness of evaluation methodologies. There is a recognized need for evaluation metrics and processes that encompass broader software quality attributes such as code maintainability, readability, and interpretability, which receive less attention currently [180].
@@
A significant challenge with the reliability of LLM generated code is the occurrence of hallucinations. Similar to issues observed in natural language generation, these models can produce code that appears plausible but deviates from the user specified requirements, contradicts contextual information, contains unnecessary repetitions or non functional dead code segments, or misuses programming interfaces and identifiers based on incorrect knowledge [181]. Studies confirm that even sophisticated models face difficulties in recognizing these hallucinations, and mitigating them effectively through prompting alone is even more challenging, highlighting a critical area for future research to ensure the trustworthiness of generated code [181]. Furthermore, the complexity of certain domains, like hardware description language generation, introduces unique hurdles [105]. Comparing different models reveals that those specifically pretrained on large code datasets might generate better formatted or more idiomatic code compared to general purpose language models, although both can demonstrate reasoning for testing logic. Output quality also shows sensitivity to the specific prompts used [132].

### v4

@@ 3.3 WLAM for Semantic Communications
SemCom is a key technology for future 6G systems, promising superior communication efficiency by directly optimizing information transfer at the semantic level [gunduz2022beyond, qin2021semantic, lan2021semantic, ZHAO2024107055, 10233741]. Large AI models are pivotal for realizing SemCom, providing the essential capability to extract and process semantic information from diverse data modalities like images, text, audio, video, and knowledge graphs, tailored to specific tasks such as classification and reconstruction. Key challenges for large AI models in SemCom include accurately identifying and precisely transmitting the desired meaning within communication systems, and ensuring robustness against channel variations through adaptive designs and re-transmission mechanisms. Leveraging the power of large AI models to address these challenges is anticipated to unlock substantial performance gains in semantic communication efficiency and reliability. The communication flow chart is demonstrated in Fig. 10.
@@ 3.3.1 Semantic Communications Preliminaries
Understanding the theoretical underpinnings of SemCom is crucial for analyzing its performance limits and guiding system design [niu2024mathematical, 9953095]. Early work reintroduced the concept of semantic entropy, positing that semantic units exhibit logical connections rather than being random [gunduz2022beyond]. Building on this, the idea of utilizing shared knowledge bases between communicating parties to exchange semantic information highlighted its potential [choi2022unified]. Subsequent research provided more formal definitions, such as considering multiple source units as equivalent if they share the same meaning (synonymous mapping), leading to generalized definitions of semantic entropy and insights into the fundamental limits of SemCom, particularly for many-to-one source scenarios [ma2023theory]. It is recognized that this synonymous mapping, where different expressions convey identical meaning (e.g., She appeared happy vs. She appeared joyful), is a key factor enabling SemCom to potentially outperform traditional bit-level communication. These developments have contributed to establishing a systematic framework for semantic information theory.
@@ 3.3.2 Modality/Task Tailored SemCom
In text domain, the authors in [farsad2018deep] first proposed deep learning based joint source and channel coding (DeepJSCC) for text transmission. They developed an recurrent neural network (RNN)-based network to directly map text to channel symbols, bypassing the separate steps of data compression and interference resistance by source coding and channel coding. This method demonstrated a lower word error rate (WER) compared to conventional methods by effectively leveraging the structural feature of text. Building on the initial idea of using AI for feature extraction, a comprehensive transmission framework for text called DeepSC was proposed in [xie2021deep]. DeepSC employs the advanced transformer architecture for text feature extraction and reconstruction. Training the transformer model in an end-to-end manner significantly improved performance. Additionally, they introduced a new performance metric called bilingual evaluation understudy (BLEU), which measures the similarity between two sentences in the feature space. Unlike WER, BLEU aligns more closely with human-like semantic definitions. To address the complexity concerns, a lite distributed semantic communication system called L-DeepSC was proposed in [xie2020lite], where network pruning and quantization are adopted to reduce the computation complexity of network part. Recognizing the importance of semantic-aware metric, the authors in [wang2022performance] further proposed a metric of semantic similarity (MSS) that compares two sentence through their graph similarity. Based on which, they optimizes the system performance by reinforcement learning with MSS as reward.
@@
In image domain, substantial research efforts have been devoted to improving SemCom. Initially, the authors in [bourtsoulatze2019deep] proposed the CNN models for image transmission, employing convolutional layers for encoding and transpose convolutional layers for decoding. Their experiments demonstrated the potential of DeepJSCC in image transmission by significantly outperforming the conventional schemes (i.e., JPEG+LDPC) in the low SNR regime. This approach was further refined in [kurka2020deepjscc]. With the advent of powerful vision transformers, researchers began exploring transformer-based DeepJSCC networks. A vanilla transformer model for DeepJSCC was proposed in [dai2022nonlinear]. This model was subsequently improved in [yang2024swinjscc, wu2023vision] by replacing the original transformer block with swin transformer block. Recently, the authors in [wu2024mambajscc] introduced the mamba model for DeepJSCC, achieving better performance than transformer-based JSCC.
@@
In the video domain, the authors in [tung2022deepwive] were the first to consider end-to-end DeepJSCC for video transmission. They classified video frames into key frames and non-key frames. Key frames are transmitted using the image-style JSCC model proposed in [xu2021wireless], while non-key frames are transmitted by encoding the residual information relative to the key frames. Additionally, they used reinforcement learning to achieve resource allocation at the frame level based on motion strength, demonstrating numerical results that overcome the cliff-effect present in conventional transmission schemes. This approach was further improved in [zhang2023deep]. Subsequently, the authors in [wang2022wireless] proposed using the last frame as a condition for the encoding and decoding process. This paradigm allows for the full exploitation of temporal dependence and enables resource allocation for each feature embedding. Their experiments showed better performance than the conventional separate source and channel coding schemes. More recently, the authors in [du2024object] considered video transmission in static scenarios, representing the movement of objects as a graph. This approach achieved better performance than conventional schemes.
@@
In audio domain, similarly, the structural feature of speech signal can be leveraged to reduce the transmission overhead or combat the channel noise. The authors in [weng2021semantic] first proposed a deep learning based SemCom system for speech signals, called DeepSC-S. They utilised the attention mechanism to identify the important information and transmit them with larger power. This approach demonstrated their effectiveness by outperforming the conventional audio transmission scheme under the common speech signals metrics. Realizing the correlation between speech and text, DeepSC-S was further extended to speech synthesis [weng2023deep]. In the speech to speech task, the authors in [han2022semantic] considered two scenarios, that is, speech recognition and reconstruction. For the recognition task, the semantic-relevant information was extracted and transmitted to the receiver. For the reconstruction task, some additional information that is irrelevant to the semantic information is also transmitted for consistent speech synthesis.
@@
In addition to the common modalities in media form, SemCom can also be applied to knowledge graphs, which are regarded as a standard and universal modality. Knowledge graphs can effectively express the attributes and their connections, thereby denoting semantic information [zzxTCOM2025, 10550151]. The authors in [zhang2023optimization] first considered adopting knowledge graphs for wireless image transmission. They extracted the knowledge graph from the image to be transmitted selectively, and transmitted the semantic information based on its importance and the available resource blocks. Leveraging on the feature of knowledge graph, they proposed a multimodal performance metric called image-to-graph semantic similarity (ISS) to measure the transmission accuracy of semantic information. Additionally, a multi-agent RL algorithm was proposed to minimize the sum of the average transmission accuracy while satisfying ISS requirement. Then, the transmission of knowledge graph in multi-user scenario was further considered in [e26050394].
@@ 3.3.3 Generative AI Models for SemCom
DeepJSCC has marked significant progress in SemCom, primarily by minimizing the distortion between the reconstructed data and the original source. However, just focusing on low distortion does not always guarantee optimal semantic performance or perceptual quality, reflecting the inherent rate-distortion-perception tradeoff [erdemir2023generative, yilmaz2023high]. To bridge this gap, researchers are increasingly integrating generative AI models into SemCom systems. These models, capable of learning complex data distributions and generating high-fidelity samples, offer powerful tools to enhance semantic fidelity and perceptual quality beyond traditional distortion metrics. Two main categories of generative models are being explored below.
@@
This line of research focuses on using established generative techniques like GANs and diffusion models to improve the quality of data reconstructed by DeepJSCC, particularly for perceptual tasks. One strategy involves employing generative models as post-processing modules. After DeepJSCC reconstructs the data, a generative model refines the output to look more realistic, such as using StyleGAN for face images [erdemir2023generative] or diffusion models for general image restoration [yilmaz2023high]. The primary challenge in wireless communications is accurately modeling the complex distortion introduced by both the DeepJSCC model and the noisy wireless channel [chen2024commin]. An alternative approach leverages the inherent denoising capabilities of diffusion models more directly, using them either for preprocessing or as integrated denoisers at the receiver to combat channel impairments [wu2024cddm]. Recognizing the computational and bandwidth demands of these models, techniques like integrating compression networks [guo2024diffusion] and knowledge distillation [pei2024latent] have been explored to improve efficiency.
@@
Leveraging the unique capabilities of LLMs opens fundamentally new avenues for SemCom, shifting focus towards deep semantic understanding and knowledge utilization. LLMs can serve as expansive knowledge bases, enabling highly efficient communication where the transmitter sends minimal, semantically crucial information, and the receiver-side LLM reconstructs the full meaning using its stored knowledge and contextual reasoning [wang2024large, ni2024interplay]. Furthermore, the training objective of LLMs, related to entropy minimization [deletang2023language], makes them inherently effective source coders capable of semantic compression. The extension of LLMs to multimodal data allows for their integration into multimodal SemCom systems, enhancing information exchange across diverse formats. Finally, there is a synergistic interplay where communication can update the knowledge of LLM, and the refined knowledge, in turn, improves communication efficiency and context-awareness.
@@ 3.4 Wireless Agents
Large AI models like LLMs are transitioning from tools for specific tasks like text generation or analysis into more autonomous entities, known as agents capable of planning reasoning and interacting with environments to achieve complex goals [xiao2025agenticainetworking6g, jiang2025largeaimodelsagentic, xi2025rise, wang2025bridgingphysicaldigitalworlds]. This agentic paradigm leverages the advanced understanding reasoning and generation capabilities inherent in LLMs applying them to orchestrate complex systems like future wireless networks. In 6G communications LLMs as agents promise to manage network complexity automate operations and enable novel intelligent services by perceiving the network state making informed decisions and interacting with network functions through defined interfaces or APIs [xiao2024llmagents6gorchestrator, dev2025advancedarchitecturesintegratedagentic], as illustrated in Fig. 11. This shift requires specific system designs model training methodologies and interaction frameworks to harness their potential effectively and reliably within the telecommunications domain. Implementing these complex agentic systems is further facilitated by the development of dedicated platforms, such as the open-source OpenManus [openmanus2025] framework designed for building general AI agents.
@@ 3.4.1 Telecom Knowledge Agent
LLMs are increasingly being adapted to function as specialized telecom knowledge agents, capable of navigating the complexities of the telecommunications domain. TelecomGPT offers a representative framework for this adaptation, proposing a comprehensive pipeline to transform general purpose LLMs into telecom specific variants [zou2024telecomgpt]. This process utilizes collected telecom data for continual pretraining, instruction dataset construction for fine tuning, and preference data for alignment tuning. The TelecomGPT framework also introduces novel benchmarks designed to evaluate critical model capabilities, including mathematical modeling, question answering, and code generation specific to the telecom context.
Specialized models enhance information retrieval from technical documents, a crucial task for telecom professionals. TeleRoBERTa, for example, provides an extractive question answering model highly effective at referencing 3GPP documents [10624786]. By adapting the RoBERTa base model and fine tuning it on telecom data, TeleRoBERTa delivers performance comparable to much larger foundation models on benchmarks like TeleQuAD, showcasing the efficiency gains possible through domain specific training even with fewer parameters. Complementing this, TelecomRAG leverages retrieval augmented generation to create a telecommunication standards assistant [telecomrag2023]. This system focuses on providing accurate, technically detailed, and verifiable responses, directly addressing the limitations of generic models regarding precision and source traceability when dealing with standard specifications.
The development of domain specific models is further supported by initiatives like the Tele-LLMs series, which provides open source LLMs adapted for telecommunications [maatouk2024telellmsseriesspecializedlarge]. These models, enhanced through continual pretraining on telecom related data, demonstrate improved performance on domain relevant tasks compared to their general purpose predecessors. Together, these frameworks and specialized models highlight significant advancements in tailoring LLMs to serve as effective and reliable knowledge agents within the telecommunications field, promising greater efficiency and accessibility for professionals navigating complex technical information.
@@ 3.4.2 Code Generation
General applications include translating natural language descriptions into executable code, providing context aware code completion suggestions during development, and performing automatic program repair to correct bugs in existing software [10403378]. Beyond these general uses, studies have shown the applicability of these models in more specialized areas relevant to communications engineering. For instance, LLMs have been successfully employed to generate complex hardware description language code, such as Verilog, for implementing advanced wireless communication algorithms like the fast Fourier transform on field programmable gate arrays. Techniques including in context learning and chain of thought prompting were utilized to navigate the intricacies of hardware specific requirements like subtask scheduling and multi step reasoning, which are less common challenges in standard software code generation [du2024powerlargelanguagemodels]. Another targeted application involves the automated generation of test scripts essential for validating telecom software systems. A proposed two stage framework first uses generative models to create synthetic yet realistic test input data based on historical network performance, and then employs a LLM to generate the actual test script code using this synthetic data combined with natural language test descriptions [10615269]. This addresses the tedious nature of manual test creation and helps cover diverse scenarios, crucial for complex telecom environments like O-RAN.
@@
While the application potential is vast, the evaluation of code generated by LLMs presents ongoing challenges. Much current evaluation work focuses primarily on functional correctness, often assessed through pass rates on standardized programming benchmarks, and on identifying security vulnerabilities [10403378, liu2024exploringevaluatinghallucinationsllmpowered]. However, research indicates a gap between the advancement of generation capabilities and the comprehensiveness of evaluation methodologies. There is a recognized need for evaluation metrics and processes that encompass broader software quality attributes such as code maintainability, readability, and interpretability, which receive less attention currently [10403378].
@@
A significant challenge with the reliability of LLM generated code is the occurrence of hallucinations. Similar to issues observed in natural language generation, these models can produce code that appears plausible but deviates from the user specified requirements, contradicts contextual information, contains unnecessary repetitions or non functional dead code segments, or misuses programming interfaces and identifiers based on incorrect knowledge [liu2024exploringevaluatinghallucinationsllmpowered]. Studies confirm that even sophisticated models face difficulties in recognizing these hallucinations, and mitigating them effectively through prompting alone is even more challenging, highlighting a critical area for future research to ensure the trustworthiness of generated code [liu2024exploringevaluatinghallucinationsllmpowered]. Furthermore, the complexity of certain domains, like hardware description language generation, introduces unique hurdles [du2024powerlargelanguagemodels]. Comparing different models reveals that those specifically pretrained on large code datasets might generate better formatted or more idiomatic code compared to general purpose language models, although both can demonstrate reasoning for testing logic. Output quality also shows sensitivity to the specific prompts[10615269].